The two stones were thrown vertically downwards from a height of 50 m towards the surface of the lake, with the distance between their throws being 1 s. Both stones hit the surface of the lake at the same time. The initial speed of the first stone is 2m / s. Determine the speed at which the second stone is thrown vertically downwards.


1
Expert's answer
2021-03-03T11:26:04-0500

"s=v_{01}t_1+gt_1^2/2\to50=2t_1+9.8t_1^2/2\to t_1\approx3\ (s)"


"t_2=3-1=2\ (s)"


"s=v_{02}t_2+gt_2^2/2\to50=v_{02}\cdot2+9.8\cdot2^2/2\to v_{02}\approx15.2\ (s)" . Answer